The vibration of which bond gives an IR absorption
Answer #1
Concepts and reason

The structure of two molecules is given which needs to be distinguished based on the different IR stretching band shown by them. IR band strength is based on the type of bond formed and the reduced mass of the bond forming species. Both the structure may show the similar bands, therefore, identify the band which is present only in one. Every bond vibration has a distinct IR frequency range.

Fundamentals

IR, that is, infrared spectroscopy is helpful in determining the type of bond and functional group present in a molecule. Vibrations produced due to stretching or bending of a band in a molecule can be detected at a specific frequency region of IR spectra. IR spectra arise in the molecule which undergoes change in dipole moment while undergoing bond stretching or bending.

IR frequency range is about 500cm16670cm1500\,{\rm{c}}{{\rm{m}}^{ - 1}} - 6670\,{\rm{c}}{{\rm{m}}^{ - 1}} , the frequency in IR is represented in the units of wave number. Consider the following relation:

υ=12πkμ\upsilon = \frac{1}{{2\pi }}\sqrt {\frac{k}{\mu }}

Here,

υ=frequencyk=bondconstant,characteristicpropertyμ=reducedmass\begin{array}{l}\\\upsilon = {\rm{frequency}}\\\\k = {\rm{bond}}\,{\rm{constant, characteristic property}}\\\\\mu = {\rm{reduced mass}}\\\end{array}

Part a

Consider the structure of both the compounds and identify the bands obtained in both as given as shown below:

Н, Н.
H,С—с—с—он
Hа
Н.С—с—о— СH3,
II
I
|о С.С Вand
С-H band
С-С Вand
о сHband
о сO band
С-O band
(0 ОH band

Assign the bond which will differentiate between the given compound as shown below:

Н, Н.
Н,С —с—с—он
На
HаС—с—0—СH,
II
I
О ОH bond
О-H bond
absent
present

Thus, vibration of OH{\rm{O}} - {\rm{H}} bond will give an IR absorption that distinguishes between the given compound.

Part b

Label the wavenumber of the OH{\rm{O}} - {\rm{H}} band as shown below:

OH=3500cm1{\rm{O}} - {\rm{H}} = 3500\,{\rm{c}}{{\rm{m}}^{ - 1}}

Ans: Part a

C-O band
O-H band
C-C Band
C-H band